11. The range of values of a for which the equation ax2+2x - 2 = 0 has real roots is:

                 A. a ≤ 1/2            B. a ≥ 1/2            C. a ≤ -1/2        D. a ≥ -1/2

 

12. if the graph of y =x2 - 8x + m cuts the x-axis in two distinct points then

               A. m < !6                      B. m > 16                  C. m ≤ 16                  D. m ≥ 16


1
Expert's answer
2021-02-24T06:58:41-0500

11.Ans:- For an equation "ax^2+bx+c=0" , "b^2-4ac" is called the discriminant and helps in determining the nature of the roots of a quadratic equation. If "b^2-4ac>0" , the roots are real and distinct. If "b^2-4ac=0" , the roots are real and equal.

Given that "ax^2+2x-2=0" has real roots.

So, "b^2-4ac\ge0" here "b=2 , c=-2"

"\Rightarrow" "{2}^2-4\times{a}\times{-2}\ge0"

"\Rightarrow" "a\ge" "-\dfrac{1}{2}"

Therefore, option (D) is correct.

12.Ans:-

condition for the quadratic equation cuts two distinct points on "x" -axis is "\dfrac{-D}{4a}\le0" where D is "b^2-4ac".

"\Rightarrow" "-\dfrac{8^2-4\times{1}\times{m}}{4\times1}\le0"


"\Rightarrow" "m\le16"

therefore option (C) is correct